2x-y+-3
x+y=9
Please Use Substitution method

Answer by checkley79(3341)

2x-y=-3 I'll assume the + should have been a = sign.
x+y=9
x=9-y
2(9-y)-y=-3
18-2y-y=-3
-3y=-3-18
-3y=-21
y=-21/-3
y=7 ans.
x+7=9
x=9-7
x=2 ans.
Proof:
2*2-7=-3
4-7=-3
-3=-3